# On the choice of the non-trainable internal weights in random feature maps

### v1

@@ Abstract.
The computationally cheap machine learning architecture of random feature maps can be viewed as a single-layer feedforward network in which the weights of the hidden layer are random but fixed and only the outer weights are learned via linear regression. The internal weights are typically chosen from a prescribed distribution. The choice of the internal weights significantly impacts the accuracy of random feature maps. We address here the task of how to best select the internal weights. In particular, we consider the forecasting problem whereby random feature maps are used to learn a one-step propagator map for a dynamical system. We provide a computationally cheap hit-and-run algorithm to select good internal weights which lead to good forecasting skill. We show that the number of good features is the main factor controlling the forecasting skill of random feature maps and acts as an effective feature dimension. Lastly, we compare random feature maps with single-layer feedforward neural networks in which the internal weights are now learned using gradient descent. We find that random feature maps have superior forecasting capabilities whilst having several orders of magnitude lower computational cost.
Pinak Mandal,***Corresponding author: pinak.mandal@sydney.edu.au 1 Georg A. Gottwald1
@@ 1. Introduction
Estimation and prediction of the state of a dynamical system evolving in time is central to our understanding of the natural world and to controlling the engineered world. Often practitioners are tasked with such problems without the knowledge of the underlying governing dynamical system. In such scenarios a popular approach is to reconstruct the dynamical model from observations of the system [31, 1, 5]. Predicting the future state of the system from these reconstructions is particularly challenging for chaotic dynamical systems. Chaotic dynamical systems cannot be accurately predicted beyond a finite time known as the predictability time due to their sensitive dependence on the initial conditions.
@@
In recent times machine learning has achieved remarkable progress in learning surrogate models for dynamical systems from given data. Recurrent networks such as Long Short-Term Memory networks [39, 41] and gated recurrent units [8] have been successfully applied in a plethora of time series prediction tasks [9, 6, 22]. These methods however often contain learnable parameters of the order of $\mathcal{O}(10^{6})$, and require substantial fine tuning of hyperparameters and costly optimization strategies [21]. An attractive alternative is provided by random feature maps [36, 37, 33] and its extensions such as echo state networks and reservoir computers [30, 29, 35, 32]. These architectures can be viewed as a single-layer feedforward network in which the weights and biases of the hidden layer, the so called internal parameters, are randomized before training and then are kept fixed. This renders the costly nonconvex optimization problem of neural networks to a simple linear least-square regression for the outer weights. The output of random feature maps and its extensions is hence a linear combination of a high-dimensional randomized basis. These methods have been shown to enjoy the universal approximation property, which states that in principle they can approximate any continuous function arbitrarily close [38, 3, 16, 12].
@@
We focus here on classical random feature maps [36, 37, 33] which have recently been shown to have excellent forecasting skill for chaotic dynamical systems [15, 14]. The fact that random feature maps enjoy the universal approximation property does not provide practitioners with information on how to choose the internal parameters. The internal parameters are typically drawn from some prescribed distribution such as the uniform distribution on an interval or a Gaussian distribution. The forecasting capability of the learned surrogate map sensitively depends on the choice of the distribution [15]. To generate good internal parameters which lead to improved performance of random feature maps, several data-independent methods such as Monte Carlo and quadrature based algorithms as well as data-dependent methods such as leverage score based sampling and kernel learning have been proposed; for a detailed survey see [23]. In recent work Dunbar et al [34] choose the distribution of the random weights from a parametric family. The parameters are chosen to optimize a cost function motivated from Empirical Bayes arguments, with the optimization performed with derivative-free Ensemble Kalman inversion. Here we introduce a computationally cheap, non-parametric, optimization-free and data-driven method to draw internal parameters which lead to improved forecasting skill. We argue that good features, corresponding to good internal parameters, need to explore the expressivity of a given activation function. We consider here as an example the $\tanh$ activation function. To allow for good expressivity, good parameters should neither map the training data into the linear range of the activation function nor into the saturated range in which different inputs cannot be discerned. This leads us to a definition of good features corresponding to good internal parameters. We show that the set of good internal parameters is non-convex but can be expressed as a union of convex sets. To sample from a convex set we employ a hit-and-run algorithm [40, 43]. Hit-and-run algorithms are a class of Markov chain Monte Carlo samplers known for their fast mixing times in convex regions [26, 28, 20]. In recent years, hit-and-run algorithms have also been analyzed for sampling nonconvex regions [7, 17, 2].
@@
A secondary objective of our work is to demonstrate that a random feature map typically achieves superior forecasting skill when compared to a neural network of the same architecture, trained with gradient descent, while being several orders of magnitude cheaper computationally. We show that the bad performance of the single-layer feedforward network can be attributed to the optimization procedure not being constrained to the set of good internal parameters. This can potentially lead to new design and improved training schemes for more complex networks.
The outline of this paper is as follows. In Section 2 we describe the setup of data-driven surrogate maps for dynamical systems and how to assess their forecasting capabilities. Section 3 introduces random feature maps and illustrates how the choice of the internal weights effects the forecasting capabilities of the associated trained surrogate maps. Section 4 defines the set of good internal parameters and introduces hit-and-run algorithms to uniformly sample from this set. Section 5 illustrates the effect of sampling from the good set of internal parameters on the forecasting skill and how the least-square training learns to distinguish good features associated with good parameters from those associated with internal parameters drawn from the complement of the good set. Section 6 compares random feature maps with single-layer feedforward networks in which the internal parameters are learned using backpropagation, and establishes that random feature maps with good parameters far outperform the single-layer feedforward neural network. Finally, we conclude in Section 7 with a summary of our results and possible future extensions.
@@ 2. Dynamical setup
We consider the forecasting problem for chaotic dynamical systems. Consider the following $D$-dimensional continuous-time system,
@@
with initial data $\mathbf{u(0)}=\mathbf{u}_{0}$, which we observe at discrete times $t_{n}=n\Delta t$ for $n=1,2,\ldots,N$. We consider here the case when the full $D$-dimensional state is observed and observations are noise-free. For the treatment of noisy observations and partial observations see [15, 14]. We view the dynamical system of these observations in terms of a discrete propagator map,
@@
The aim of data-driven modelling is to construct a surrogate map $\Psi_{S}$ from the training data given by the observations that well approximates the true propagator map $\Psi_{\Delta t}$ of (2). In the following we denote variables associated with the surrogate map with a hat, and write the learned surrogate dynamical system as
$\hat{\mathbf{u}}_{n+1}=\Psi_{S}(\hat{\mathbf{u}}_{n}),$ (3)
@@
as the underlying continuous dynamical system (1). The Lorenz-63 system is chaotic with a positive Lyapunov exponent of $\lambda_{\rm{max}}\approx 0.91$ [42]. We generate independent training and validation data sampled at $\Delta t=0.02$ by randomly selecting initial conditions $\mathbf{u}_{0}$. We discard an initial transient dynamics of $40$ time units to ensure that the dynamics has settled on the attractor.
To test the predictive capability of a surrogate model, we define the forecast time $\tau_{f}$ associated with the surrogate model,
@@ 3. Random feature maps
$\Psi_{S}(\mathbf{u})=\mathbf{W}\sigma{(\mathbf{W}_{\rm in}\mathbf{u}+\mathbf{b_{\rm in}})},$ (6)
where $\mathbf{u}$ is the $D$-dimensional state vector, $\mathbf{W}_{\rm in}\in\mathbb{R}^{D_{r}\times D}$ is the internal weight matrix, $\mathbf{b}_{\rm in}\in\mathbb{R}^{D_{r}}$ the internal bias and $\mathbf{W}\in\mathbb{R}^{D\times D_{r}}$ the outer weight matrix. The nonlinear activation function $\sigma$ is applied component wise and we choose here $\sigma=\tanh$. Random features are characterized by the internal weights $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ being drawn before training from a prescribed distribution $p(w_{\rm in})$ and $p(b_{\rm in})$. The internal weights remain fixed and are not learned as it would be the case for a single-layer feedforward network which has the same architecture as in (6). Random feature maps can hence be seen as a linear combination of $D_{r}$-dimensional random features vectors
@@
In the following we refer to components of this feature vector as features.
The matrix $\mathbf{W}$, controlling the linear combinations of the feature vectors (7), is learned from training data $\mathbf{U}\in\mathbb{R}^{D\times N}$ the columns of which are the observations $\mathbf{u}_{n},\;n=1,\ldots,N$ of the system (2). We do so by solving the following regularized optimization problem,
$\mathbf{W}^{*}=\underset{\mathbf{W}}{\operatorname*{arg\,min}}\;\mathcal{L}(\mathbf{W};\mathbf{U}),$ (8)
@@
$\mathcal{L}(\mathbf{W};\mathbf{U})=\|\mathbf{W}\mathbf{\Phi}(\mathbf{U})-\mathbf{U}\|^{2}+\beta\|\mathbf{W}\|^{2}.$ (9)
@@
The low computational cost of random feature maps makes them a very attractive architecture.
@@ 3.1. The effect of the internal weights on the performance of random feature maps
Random feature maps enjoy the universal approximation property [38, 3] and hence, in principle, for a sufficiently high feature dimension $D_{r}$ can approximate continuous functions arbitrarily close. The universal approximation property however does not guide practitioners how to find the internal weights $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ which allow for such an approximation. The main objective of this paper is to sample the internal parameters in a way that increases the forecasting skill of the random feature maps.
@@
Indeed, the forecasting skill of a learned random feature map surrogate model (3) sensitively depends on the internal weights. To illustrate the effect of the hyperparameters $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ on the forecast time $\tau_{f}$ we uniformly sample $\mathbf{W}_{\rm in}$ and $\mathbf{b}_{\rm in}$ from the intervals $[-w,w]$ and $[-b,b]$, respectively, with $(w,b)\in(0,w_{\rm max})\times(0,b_{\rm max})$. In particular, we use $30\times 30$ regular grid points over $(0,w_{\rm max})\times(0,b_{\rm max})$ with $w_{\rm max}=0.4$ and $b_{\rm max}=4.0$, and probe the statistics by generating $M=100$ feature maps for each grid point, while keeping the training data and the validation data fixed for all realizations to focus on the effect of the internal weights. We fix the feature dimension at $D_{r}=300$ and the regularization parameter at $\beta=4\times 10^{-5}$. Figure 1 shows a contour plot of the mean and the standard deviation of the forecast time $\tau_{f}$ over the domain of the internal weights. We can clearly see that certain regions in the hyperparameter space are associated with good performance with mean forecast times $\tau_{f}>4$ while other regions produce poor mean forecast times. Moreover, regions in the hyperparameter space corresponding to high mean forecast times $\tau_{f}$ may have large variance.
@@
Figure 1. Contour plots of the mean and standard deviation of the forecast time $\tau_{f}$ computed using $\mathbf{W}_{\rm in},\mathbf{b}_{\rm in}$ sampled uniformly from intervals of variable size $[-w,w]$ and $[-b,b]$ respectively. Samples were drawn for grid points $(w,b)$ on a $30\times 30$ regular grid over the domain $(0,0.4)\times(0,4.0)$. Averages are taken over $M=100$ realizations per grid-point $(w,b)$, for a feature dimension $D_{r}=300$, training data length $N=20,000$ and regularization parameter $\beta=4\times 10^{-5}$, using fixed training and validation data.
@@
Ideally, we would like parameters which have both, high mean forecast time and low variance so that the performance is not dependent on the particular training data used. It is clear that if the internal weights $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ are chosen sufficiently small, the associated features (7) are essentially linear with $\boldsymbol{\phi}\approx\mathbf{W}_{\rm in}\mathbf{u}+\mathbf{b_{\rm in}}$ for all input data $\mathbf{u}$. This would render the random feature maps a linear model which are known to be incapable of modelling nonlinear chaotic dynamical systems [10, 4]. On the other extreme, for sufficiently large internal weights a $\tanh$-activation function saturates, and one obtains $\boldsymbol{\phi}\approx\pm 1$ independent of the input data $\mathbf{u}$, severely decreasing the expressivity of the random feature map. This suggests that one should choose internal weights which sample the $\tanh$-activation function in its nonlinear non-saturated range. This is illustrated in Figure 2. We shall call features linear, if for all data $\mathbf{u}$ the argument of the $\tanh$-activation function lies within the interval centred around the origin in $[-L_{0},L_{0}]$. Those features obtained by the $\tanh$-activation function that are approximately ±$\pm 1$ for all input data $\bf{u}$, i.e. where the arguments of the $\tanh$-activation function lie in either of two unbounded sets $(-\infty,-L_{1}],[L_{1},+\infty)$, we label saturated features. Those features which for all input data are neither linear nor saturated, i.e. for which the argument of the $\tanh$-activation function lies in either of the two intervals $(-L_{1},-L_{0})$ or $(L_{0},L_{1})$, are labelled good features. We use $L_{0}=0.4$ and $L_{1}=3.5$ to define good, linear and saturated features throughout this paper.
@@
Figure 2. Domain and range of features produced by a $\tanh$-activation function with $L_{0}=0.4$ and $L_{1}=3.5$, leading to linear, saturated and good features respectively.
@@
To illustrate the detrimental effect of saturated features on the forecasting skill of random features we select from the random feature maps which were sampled in Figure 1, those if they fall into two groups, those that lead to particularly large forecast times $\tau_{f}>8$ and those that lead to particularly low forecast times $\tau_{f}<0.5$. For each of those feature vectors we determine the average fraction $F_{s}$ of how much of the data input ${\bf{u}}$ is mapped to the saturated values ±$\pm 1$, by probing for $800$ randomly selected data points ${\bf{u}}_{n}$. For each group we randomly select $500$ samples from the $90,000$ random feature maps used in Figure 1. Figure 3 shows that the histograms of $F_{s}$ for these two groups are clearly distinct. The group with low forecast times $\tau_{f}<0.5$ has a significantly higher probability of having more saturated features compared to the group with large forecast times $\tau_{f}>8$. The pronounced peak at $F_{s}=0$ is a sampling effect: when sampling uniformly from the grid $(0,w_{\rm max})\times(0,b_{\rm max})$ with $w_{\rm max}=0.4$ and $b_{\rm max}=4.0$, it is much more likely to draw parameters which correspond to non-saturated features. Such random feature map samples are much more likely to have higher forecast times and hence are concentrated entirely in the $\tau_{f}>8$ group.
@@
In the following Section we develop a computationally cheap algorithm to sample from the set of good weights and show in Section 5 how this increases the forecasting skill of random feature surrogate maps.
@@
Figure 3. Empirical histograms of average fraction of saturated features $F_{s}$ for random feature maps corresponding to large forecast times with $\tau_{f}>8$ and to low forecast times with $\tau_{f}<0.5$. Each group has $500$ samples and the histograms depict the probability of having a certain value of $F_{s}$ in each group.

### v2

@@ Abstract.
The computationally cheap machine learning architecture of random feature maps can be viewed as a single-layer feedforward network in which the weights of the hidden layer are random but fixed and only the outer weights are learned via linear regression. The internal weights are typically chosen from a prescribed distribution. The choice of the internal weights significantly impacts the accuracy of random feature maps. We address here the task of how to best select the internal weights. In particular, we consider the forecasting problem where random feature maps are used to learn a one-step propagator map for a dynamical system. We provide a computationally cheap hit-and-run algorithm to select good internal weights which lead to good forecasting skill. We show that the number of good features is the main factor controlling the forecasting skill of random feature maps and acts as an effective feature dimension. Lastly, we compare random feature maps with single-layer feedforward neural networks in which the internal weights are now learned using gradient descent. We find that random feature maps have superior forecasting capabilities whilst having several orders of magnitude lower computational cost.
Pinak Mandal,***pinak.mandal@sydney.edu.au, corresponding author 1 Georg A. Gottwald, georg.gottwald@sydney.edu.au 1 Nicholas Cranch, ncra8185@uni.sydney.edu.au 1
@@ 1. Introduction
Estimation and prediction of the state of a dynamical system evolving in time is central to our understanding of the natural world and to controlling the engineered world. Often practitioners are tasked with such problems without the knowledge of the underlying governing dynamical system. In such scenarios a popular approach is to reconstruct the dynamical model from observations of the system [38, 1, 46, 5]. Predicting the future state of the system from these reconstructions is particularly challenging for chaotic dynamical systems. Chaotic dynamical systems cannot be accurately predicted beyond a finite time known as the predictability time due to their sensitive dependence on the initial conditions.
@@
In recent times machine learning has achieved remarkable progress in learning surrogate models for dynamical systems from given data. Recurrent networks such as Long Short-Term Memory networks [45, 48] and gated recurrent units [8] have been successfully applied in a plethora of time series prediction tasks [10, 6, 28]. These methods however often contain learnable parameters of the order of $\mathcal{O}(10^{6})$, and require substantial fine tuning of hyperparameters and costly optimization strategies [27]. An attractive alternative is provided by random feature maps [43, 42, 40] and their extensions such as echo state networks and reservoir computers [36, 35, 41, 39]. These architectures can be viewed as a single-layer feedforward network in which the weights and biases of the hidden layer, the so-called internal parameters, are randomly selected before training and then are kept fixed. This renders the costly nonconvex optimization problem of neural networks to a simple linear least-square regression for the outer weights. The output of random feature maps and their extensions is hence a linear combination of a high-dimensional randomized basis. These methods have been shown to enjoy the universal approximation property, which states that in principle they can approximate any continuous function arbitrarily close [44, 3, 17, 13].
@@
We focus here on classical random feature maps [43, 42, 40] which have recently been shown to have excellent forecasting skill for chaotic dynamical systems [16, 15]. The fact that random feature maps enjoy the universal approximation property does not provide practitioners with information on how to choose the internal parameters. The internal parameters are typically drawn from some prescribed distribution such as the uniform distribution on an interval or a Gaussian distribution. The forecasting capability of the learned surrogate map sensitively depends on the choice of the distribution [16]. To generate good internal parameters which lead to improved performance of random feature maps, several data-independent methods such as Monte Carlo and quadrature based algorithms as well as data-dependent methods such as leverage score based sampling and kernel learning have been proposed; for a detailed survey see [29]. Levine and Stuart [26] performed Bayesian optimization to determine the internal weights together with other hyperparamters such as the regularization parameter. Dunbar et al [9] choose the distribution of the random weights from a parametric family. Its parameters are chosen to optimize a cost function motivated from Empirical Bayes arguments, with the optimization performed with derivative-free Ensemble Kalman inversion. Here we introduce a computationally cheap, non-parametric, optimization-free and data-driven method to draw internal parameters which lead to improved forecasting skill. We argue that good features, corresponding to good internal parameters, need to explore the expressive range of a given activation function. We consider here as an example the $\tanh$ activation function. To allow for good expressivity, good parameters should neither map the training data into the linear range of the activation function nor into the saturated range in which different inputs cannot be discerned. This leads us to a definition of good features corresponding to good internal parameters. We find that the set of good internal parameters is non-convex but can be expressed as a union of convex sets. To sample from a convex set we employ a hit-and-run algorithm [47, 50]. Hit-and-run algorithms are a class of Markov chain Monte Carlo samplers known for their fast mixing times in convex regions [32, 34, 25]. In recent years, hit-and-run algorithms have also been analyzed for sampling nonconvex regions [7, 22, 2].
@@
A secondary objective of our work is to demonstrate that a random feature map typically achieves superior forecasting skill when compared to a neural network of the same architecture, trained with gradient descent, while being several orders of magnitude cheaper computationally. We show that the bad performance of the single-layer feedforward network can be attributed to the optimization procedure not being constrained to the set of good internal parameters. This can potentially lead to new design and improved training schemes for more complex network architectures.
The outline of this paper is as follows. In Section 2 we describe the setup of data-driven surrogate maps for dynamical systems and how to assess their forecasting capabilities. Section 3 introduces random feature maps and illustrates how the choice of the internal weights affects the forecasting capabilities of the associated trained surrogate maps. Section 4 defines the set of good internal parameters and introduces hit-and-run algorithms to uniformly sample from this set. Section 5 illustrates the effect of sampling from the good set of internal parameters on the forecasting skill and how the least-square training learns to distinguish good features associated with good parameters from those associated with internal parameters drawn from the complement of the good set. Section 6 compares random feature maps with single-layer feedforward networks in which the internal parameters are learned using backpropagation, and establishes that random feature maps with good parameters far outperform the single-layer feedforward neural network. Section 7 shows that random feature maps with good internal parameters reproduce the long-time statistical behaviour of the underlying dynamical system more accurately than those without. Finally, we conclude in Section 8 with a summary of our results and possible future extensions.
@@ 2. Dynamical setup
We consider the forecasting problem for chaotic dynamical systems. Consider the following $D$-dimensional continuous-time dynamical system,
@@
with initial data $\mathbf{u(0)}=\mathbf{u}_{0}$, which we observe at discrete times $t_{n}=n\Delta t$ for $n=0,1,\ldots,N$. We consider here the case when the full $D$-dimensional state is observed and observations are noise-free. For the treatment of noisy observations and partial observations see [16, 15]. We view the dynamical system of these observations in terms of a discrete propagator map,
@@
The aim of data-driven modelling is to construct a surrogate map ${\hat{\Psi}}_{\Delta t}$ from the training data given by the observations that well approximates the true propagator map $\Psi_{\Delta t}$ of (2). In the following we denote variables associated with the surrogate map with a hat, and write the learned surrogate dynamical system as
$\hat{\mathbf{u}}_{n+1}={\hat{\Psi}}_{\Delta t}(\hat{\mathbf{u}}_{n}),$ (3)
@@
as the underlying continuous dynamical system (1). The Lorenz-63 system is chaotic with a positive Lyapunov exponent of $\lambda_{\rm{max}}\approx 0.91$ [49]. We generate independent training and validation data, $\mathbf{u}_{n}$ and $\mathbf{u}_{n}^{\rm validation}$, sampled at $\Delta t=0.02$ by randomly selecting independent initial conditions $\mathbf{u}_{0}$ and $\mathbf{u}_{0}^{\rm validation}$, respectively. We discard an initial transient dynamics of $40$ time units to ensure that the dynamics has settled on the attractor.
To test the predictive capability of a surrogate model, we apply it to unseen validation data and define the forecast time $\tau_{f}$ associated with the surrogate model,
@@ 3. Random feature maps
${\hat{\Psi}}_{\Delta t}(\mathbf{u})=\mathbf{W}\sigma{(\mathbf{W}_{\rm in}\mathbf{u}+\mathbf{b_{\rm in}})},$ (6)
where $\mathbf{u}$ is the $D$-dimensional state vector, $\mathbf{W}_{\rm in}\in\mathbb{R}^{D_{r}\times D}$ is the internal weight matrix, $\mathbf{b}_{\rm in}\in\mathbb{R}^{D_{r}}$ the internal bias and $\mathbf{W}\in\mathbb{R}^{D\times D_{r}}$ the outer weight matrix. The nonlinear activation function $\sigma$ is applied component wise and we choose here $\sigma=\tanh$. Random features are characterized by the internal weights $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ being drawn before training from a prescribed distribution $p(w_{\rm in},b_{\rm in})$. The internal weights remain fixed and are not learned as it would be the case for a single-layer feedforward network which has the same architecture as in (6). Random feature maps can hence be seen as a linear combination of features which are the components of the $D_{r}$-dimensional random features vectors
@@
The matrix $\mathbf{W}$, controlling the linear combinations of the features, is learned from training data $\mathbf{U}\in\mathbb{R}^{D\times N}$, the columns of which are the observations $\mathbf{u}_{n},\;n=1,\ldots,N$ of the system (2). We do so by solving the following regularized optimization problem,
$\mathbf{W}^{*}=\underset{\mathbf{W}}{\operatorname*{arg\,min}}\;\mathcal{L}(\mathbf{W};\mathbf{U},{\color[rgb]{0,0,0}\mathbf{W}_{\rm in},\mathbf{b_{\rm in}}}),$ (8)
@@
$\mathcal{L}(\mathbf{W};\mathbf{U},{\color[rgb]{0,0,0}\mathbf{W}_{\rm in},\mathbf{b_{\rm in}}})=\|\mathbf{W}\mathbf{\Phi}(\mathbf{U})-\mathbf{U}\|^{2}+\beta\|\mathbf{W}\|^{2}.$ (9)
@@
The low training cost of random feature maps makes them a very attractive architecture. Training reduces to the determination of the $D_{r}D$ outer weights $\mathbf{W}^{*}$ by means of an explicit analytical formula once the inner weights have been set to their fixed values. The total number of parameters for random feature maps is $D_{r}(2D+1)$ which scales linearly in the feature dimension $D_{r}$ unlike other architectures such as echo state networks [20, 21] and long short term memory (LSTM) networks [19].
@@ 3.1. The effect of the internal weights on the performance of random feature maps
Random feature maps enjoy the universal approximation property [44, 3]. This means that for a given feature dimension $D_{r}$ there exist internal weights such that the random feature map approximates any continuous function. The approximation can get arbitrarily close for increasing feature dimension. The universal approximation property, however, does not guide practitioners on how to find the internal weights $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ which allow for such an approximation. The main objective of this paper is to sample the internal parameters in a way that increases the forecasting skill of the random feature maps when compared to the usually employed data-uninformed random draws from a specified distribution such as a Gaussian or a uniform distribution.
@@
Indeed, the forecasting skill of a learned random feature map (3) sensitively depends on the internal weights. To illustrate the effect of the hyperparameters $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ on the forecast time $\tau_{f}$ we uniformly sample $\mathbf{W}_{\rm in}$ and $\mathbf{b}_{\rm in}$ from the intervals $[-w,w]$ and $[-b,b]$, respectively, with $(w,b)\in(0,w_{\rm max})\times(0,b_{\rm max})$. In particular, we use $30\times 30$ regular grid points over $(0,w_{\rm max})\times(0,b_{\rm max})$ with $w_{\rm max}=0.4$ and $b_{\rm max}=4.0$, and probe the statistics by generating $M=100$ feature maps for each grid point, while keeping the training data and the validation data fixed for all realizations to focus on the effect of the internal weights. We fix the feature dimension at $D_{r}=300$ and the regularization parameter at $\beta=4\times 10^{-5}$. Figure 1 shows a contour plot of the mean and the standard deviation of the forecast time $\tau_{f}$ over the domain of the internal weights. We can clearly see that certain regions in the hyperparameter space are associated with good performance with mean forecast times $\tau_{f}>4$ while other regions produce poor mean forecast times. Moreover, regions in the hyperparameter space corresponding to high mean forecast times $\tau_{f}$ may have large variance.
@@
Figure 1. Contour plots of the mean and standard deviation of the forecast time $\tau_{f}$ computed using $\mathbf{W}_{\rm in},\mathbf{b}_{\rm in}$ sampled uniformly from intervals of variable size $[-w,w]$ and $[-b,b]$ respectively. Samples were drawn for grid points $(w,b)$ on a $30\times 30$ regular grid over the domain $(0,0.4)\times(0,4.0)$. Averages are taken over $M=100$ realizations per grid-point $(w,b)$, for a feature dimension $D_{r}=300$, training data length $N=20,000$ and regularization parameter $\beta=4\times 10^{-5}$, using the same training and validation data for ecah realization.
@@
Ideally, we would like parameters which have both, high mean forecast time and low variance so that the performance is not dependent on the particular training data used. It is clear that if the internal weights $(\mathbf{W}_{\rm in},\mathbf{b}_{\rm in})$ are chosen sufficiently small, the associated features (7) are essentially linear with $\boldsymbol{\phi}\approx\mathbf{W}_{\rm in}\mathbf{u}+\mathbf{b_{\rm in}}$ for all input data $\mathbf{u}$. This would reduce the random feature maps to linear models which are known to be incapable of modelling nonlinear chaotic dynamical systems [11, 4]. On the other extreme, for sufficiently large internal weights a $\tanh$-activation function saturates, and one obtains $\boldsymbol{\phi}\approx\pm 1$ independent of the input data $\mathbf{u}$, severely decreasing the expressivity of the random feature map. This suggests that one should choose internal weights which sample the $\tanh$-activation function in its nonlinear non-saturated range. This is illustrated in Figure 2. We shall call features linear, if for all data $\mathbf{u}$ the argument of the $\tanh$-activation function lies within the interval centred around the origin in $[-L_{0},L_{0}]$. Those features obtained by the $\tanh$-activation function that are approximately ±$\pm 1$ for all input data $\bf{u}$, i.e. where the arguments of the $\tanh$-activation function lie in either of two unbounded sets $(-\infty,-L_{1}],[L_{1},+\infty)$, we label saturated features. Those features which for all input data are neither linear nor saturated, i.e. for which the argument of the $\tanh$-activation function lies in either of the two intervals $(-L_{1},-L_{0})$ or $(L_{0},L_{1})$, are labelled good features. We use $L_{0}=0.4$ and $L_{1}=3.5$ to define good, linear and saturated features throughout this paper.
@@
Figure 2. Domain and range of features produced by a $\tanh$-activation function with $L_{0}=0.4$ and $L_{1}=3.5$, leading to linear, saturated and good features.
@@
We now illustrate the detrimental effect of linear and saturated features on the forecasting skill of random features. From the random feature maps that were used in Figure 1, we select those that lead to particularly large forecast times $\tau_{f}>8$ and those that lead to particularly low forecast times $\tau_{f}<0.5$. For each of those feature maps we determine the fraction of features that correspond to (a) the linear region $[-L_{0},L_{0}]$, (b) the saturated region $(-\infty,-L_{1}]\cup[L_{1},+\infty)$, and (c) the good region $(-L_{1},-L_{0})\cup(L_{0},L_{1})$, averaged over input data $\mathbf{u}_{n}$. We denote these fractions by $p_{l}$, $p_{s}$ and $p_{g}$, respectively. We consider $5,000$ randomly selected data points ${\bf{u}}_{n}$ on the attractor of the Lorenz-63 system to estimate these fractions. For each group we randomly select $500$ samples from the $90,000$ random feature maps used in Figure 1. In Figure 3 we show the histograms of the fractions $p_{g}$, $p_{l}$ and $p_{s}$ for these two groups. The group with low forecast times $\tau_{f}<0.5$ contains a significantly higher fraction of linear or saturated features compared to the group with large forecast times $\tau_{f}>8$. Conversely, the group with large forecast times $\tau_{f}>8$ contains a significantly higher fraction of good features compared to the group with low forecast times $\tau_{f}<0.5$. This confirms our hypothesis that good forecast skill is associated with an abundance of good features and a lack of linear and saturated features. We remark that the pronounced peak at $p_{s}=0$ is a sampling effect: when sampling uniformly from the grid $(0,w_{\rm max})\times(0,b_{\rm max})$ with $w_{\rm max}=0.4$ and $b_{\rm max}=4.0$, it is much more likely to draw parameters which correspond to non-saturated features. Such random feature map samples are much more likely to have higher forecast times and hence are concentrated entirely in the $\tau_{f}>8$ group.
Figure 4 shows the mean forecast time $\mathbb{E}[\tau_{f}]$ as a function of the fractions of good, linear and saturated features, for the same random feature maps as shown in Figure 3. It is clearly seen that the forecast skill, as measured by the mean forecast time, monotonically improves for increasing number of good features. Similarly, the forecast skill monotonically degrades with increasing number of saturated features. The effect of the linear features is less obvious for the uniformly initialized random feature maps of Figure 1. However, as we will see later (cf. Figure 12), linear and saturated features have the same negative effect on the mean forecast time.
@@
In the following Section we develop a computationally cheap algorithm to sample from the set of good weights, and show in Section 5 how this increases the forecasting skill of random feature maps.
@@
Figure 3. Empirical histograms of average fractions of good, linear and saturated features, $p_{g}$, $p_{l}$ and $p_{s}$, respectively, for random feature maps corresponding to two groups: large forecast times with $\tau_{f}>8$ and low forecast times with $\tau_{f}<0.5$. The random feature maps are the same as those used in Figure 1. Each group contains $500$ samples and the histograms depict the probability of having a certain value of the respective fractions in each group.
Figure 4. Mean forecast time $\mathbb{E}[\tau_{f}]$ as a function of the fraction $p$ of good, linear and saturated features, respectively. The random feature maps are the same as those used in Figure 1. The mean forecast times are computed over bins $[p-\Delta p,p+\Delta p]$ with $\Delta p=0.001$. The shaded region delineates one standard deviation from the mean. We only report on bins which contain more than $100$ samples.